Ben is three times as old as his sister was two years ago. In 1 year from now Ben will be one less than twice his sister's age. How old are they now?
Answer by Paul(988)

Her sister's age 2 years ago = x
Ben's age = 3(x-2)

One year from now:
x+1= (her sister)
Ben's 3(x-2)+1

2(x+1)-1=3(x-2)+1
2x+2-1=3x-6+1
2x+1=3x-5
-x=-6
x=6

3(6-2)
3(4)=12

Hence, Ben is 12 years old and her sister is 6 years old.